OpenStudy (tmagloire1):

The position function of a particle in rectilinear motion is given by s(t) s(t) = t3 – 9t2 + 24t + 1 for t ≥ 0. Find the position and acceleration of the particle at the instant the when the particle reverses direction. Include units in your answer.

OpenStudy (mrnood):

since there are no units in the question I don't see how you can put units in the answer...

OpenStudy (tmagloire1):

My teacher said to ignore that part

OpenStudy (tmagloire1):

@misty1212

OpenStudy (irishboy123):

In order for the particle to reverse direction it must first slow to v=0. Find an equation for v(t) by differentiating s(t) and then set that to 0.

OpenStudy (tmagloire1):

@irishboy123 s(t) = t^3 – 9t^2 + 24t + 1 s'(t) = 3t^2 - 18t + 24 3t^2 - 18t + 24 = 0 3t^2 - 18t = -24 t(3t-18) = -24 t= -24 and -2

OpenStudy (tmagloire1):

@triciaal please help i really need it D:

OpenStudy (triciaal):

to check when the direction changes you need to identify turning points, the maximum and minimums

OpenStudy (tmagloire1):

Are -24 and -2 turning points?

OpenStudy (triciaal):

to get the acceleration take the second derivative of the original function

OpenStudy (tmagloire1):

so acceleration is s'(t) = 6t - 18

OpenStudy (tmagloire1):

s''(t)

OpenStudy (triciaal):